# Policy Mirror Descent with Lookahead

### v2

@@ B.1 Details and Precise Notation for Lookahead Computation
We apply again the Three-Point Descent Lemma (see e.g. Lemma A.1 in Johnson et al. [2023]) to obtain:
@@ B.3.1 Estimation Error Bound for $\hat{Q}_{h}$
Finally, each term $\mathcal{E}_{j}$ for $j\leq k$ can be bounded using the following lemma below. The proof relies on a union bound over the states reached during the construction of the search tree, see the lemma towards the end of Szepesvári [2023] for a detailed explanation and a proof.
@@ C.1 Inexact $h$-PMD with Linear Function Approximation
For any policy $\pi$ we can use the same function approximation procedure as described in our paper (see Appendix C.1) to yield a $\hat{\theta}\in\mathbb{R}^{d}$ with $\lVert Q^{\pi}-\Psi\hat{\theta}\rVert_{\infty}\leq\epsilon_{0}$, (for some value of $\epsilon_{0}$ to be defined below) by simply plugging in $h=1$. Note that this particular case was previously analysed in Szepesvári [2023] (Lecture 8). We reuse an intermediate result therein (equation (7)) which guarantees the bound:
@@
We would like to add an additional comment regarding possible ways to relax Assumption 6.2 even for the standard case where $h=1$. Recently, Mei et al. Mei et al. [2024] proposed an interesting alternative approach relying on order-based conditions (instead of approximation error based ones) to analyze policy gradient methods with linear function approximation. We would like to highlight though that their work only applies to the bandit setting to the best of our knowledge. The extension to MDPs is far from being straightforward and would require some specific investigation. It would be interesting to see if one can relax Assumption 6.2 using similar ideas.
@@ C.3 Convergence Analysis of Inexact $h$-PMD using Function Approximation
The proof of this theorem is identical to the one in Lattimore and Szepesvari [2019], up to using our new targets and error propagation analysis instead of the typical function approximation targets for $Q$-function estimation (without lookahead). The subset $\mathcal{C}\subseteq\mathcal{S}\times\mathcal{A}$ and distribution $\rho:\mathcal{C}\to\left[0,1\right]$ are given by the Kiefer-Wolfowitz theorem stated above. We suppose we have access to such a set. Note that this assumption is key to controlling the extrapolation error of a least squares estimate $\Psi\hat{\theta}$. ∎
@@ D.1 Simulations for $h$-PMD with Euclidean Regularization
Here we provide details of an additional experiment running $h$-PMD with Euclidean regularization. As in section 7, experiments were run on the bsuite DeepSea environment Osband et al. [2020] with $\gamma=0.99$ and a grid size of 64x64. The figures are identical in structure to the ones in 7, 32 runs were performed for each value of $h$ in 1, 5, 10, 15, and 20. In Fig. 3, we report the number of samples to the environment at each iteration for the same runs. When compared to the figures showing the convergence, it is clearly seen that running the algorithm with higher values of lookahead results in convergence using fewer overall samples.
@@ D.4 About the Stepsizes in Simulations
In our experiments we use the exact same stepsizes described in our theory. We also observe in our simulations that the stepsize does not go to infinity but rather decreases at some point and/or stabilizes. This is a consequence of the adaptivity of the stepsize discussed above. We also note that the simulations reported in Johnson et al. [2023] (appendix D p. 18) also validate the fact that the stepsizes do not go to infinity (see the right plots therein for the stepsizes). We also observe a similar behavior for the stepsizes in our experiments (see Figure 5).
@@ D.6 Experiments with MCTS-based Lookahead Estimation
We have performed additional experiments in an even more challenging and practical setting, on two environments from the bsuite Osband et al. [2020] set: DeepSea and Catch. We have implemented our PMD algorithm with a Monte Carlo Tree Search (MCTS) algorithm (which we call PMD-MCTS) to compute the lookahead Q-function. We used DeepMind’s MCTS implementation (MCTX) [^3] https://github.com/google-deepmind/mctx in order to run this algorithm on any deterministic gym style environments implemented in JAX (we use gymnax Lange [2022] to implement these environments). Note that any tree search algorithm implementation can be used to estimate the lookahead function and plugged into our PMD method. See figures below. Note that this training procedure is more practical: we do not evaluate the value function at each state-action pair anymore at each time step, we only evaluate the value function (and so only update the policy) in states that we encounter using the current policy. This significantly relaxes the generative model assumption. Note that in this more challenging setting the best performance is not obtained for higher values of $h$: intermediate values of $h$ perform better, illustrating the tradeoff in choosing the depth $h$.
@@ D.7 Continuous Control Experiments
Finally, we performed experiments in classic continuous control environments in order to investigate the performance of $h$-PMD when applied to MDPs with large/infinite state spaces. For this continuous state space setting, we implemented a fully online variant of $h$-PMD based off of DeepMind’s MuZero Schrittwieser et al. [2019] algorithm. Experiments show that increasing the lookahead depth can lead to better performance up to a certain point, and the tradeoff seems to be environment-specific.
Figure 8: Performance of $h$-PMD in two of OpenAI’s gym Brockman et al. [2016] environments, CartPole-v1 (left) and Acrobot-v1 (right). In both figures the reward is plotted against the total number of environment steps evaluated, which can be understood as the number of samples used. Note that higher values of $h$ do lead to better performance for the CartPole-v1 environment, but not necessarily for the Acrobot-v1 environment.
@@ D.8 Additional experiments
We conclude with results from a larger set of experiments designed to verify the robustness of our experimental results to different environment parameters. We report the performance of the $h$-PMD algorithm in the inexact setting in two different environments, DeepSea from bsuite Osband et al. [2020], and the classic Gridworld environment. The effect of $\gamma$ is as expected: higher values of $\gamma$ result in slower convergence, lower values of $\gamma$ yield faster convergence. The size of the environment (“chain length” for DeepSea and number of columns/rows for Gridworld) directly affects the number of iterations until convergence predictably: larger environments take more iterations to solve. Note that in all cases the effect of $h$ is as expected, i.e. higher $h$ leads to faster convergence.

### v3

@@ B.1 Details and Precise Notation for Lookahead Computation
We apply again the Three-Point Descent Lemma (see e.g. Lemma A.1 in [Johnson et al., 2023]) to obtain:
@@ B.3.1 Estimation Error Bound for $\hat{Q}_{h}$
Finally, each term $\mathcal{E}_{j}$ for $j\leq k$ can be bounded using the following lemma below. The proof relies on a union bound over the states reached during the construction of the search tree, see the lemma towards the end of [Szepesvári, 2023] for a detailed explanation and a proof.
@@ C.1 Inexact $h$-PMD with Linear Function Approximation
For any policy $\pi$ we can use the same function approximation procedure as described in our paper (see Appendix C.1) to yield a $\hat{\theta}\in\mathbb{R}^{d}$ with $\lVert Q^{\pi}-\Psi\hat{\theta}\rVert_{\infty}\leq\epsilon_{0}$, (for some value of $\epsilon_{0}$ to be defined below) by simply plugging in $h=1$. Note that this particular case was previously analysed in [Szepesvári, 2023] (Lecture 8). We reuse an intermediate result therein (equation (7)) which guarantees the bound:
@@
We would like to add an additional comment regarding possible ways to relax Assumption 6.2 even for the standard case where $h=1$. Recently, Mei et al. [Mei et al., 2024] proposed an interesting alternative approach relying on order-based conditions (instead of approximation error based ones) to analyze policy gradient methods with linear function approximation. We would like to highlight though that their work only applies to the bandit setting to the best of our knowledge. The extension to MDPs is far from being straightforward and would require some specific investigation. It would be interesting to see if one can relax Assumption 6.2 using similar ideas.
@@ C.3 Convergence Analysis of Inexact $h$-PMD using Function Approximation
The proof of this theorem is identical to the one in [Lattimore and Szepesvari, 2019], up to using our new targets and error propagation analysis instead of the typical function approximation targets for $Q$-function estimation (without lookahead). The subset $\mathcal{C}\subseteq\mathcal{S}\times\mathcal{A}$ and distribution $\rho:\mathcal{C}\to\left[0,1\right]$ are given by the Kiefer-Wolfowitz theorem stated above. We suppose we have access to such a set. Note that this assumption is key to controlling the extrapolation error of a least squares estimate $\Psi\hat{\theta}$. ∎
@@ D.1 Simulations for $h$-PMD with Euclidean Regularization
Here we provide details of an additional experiment running $h$-PMD with Euclidean regularization. As in section 7, experiments were run on the bsuite DeepSea environment [Osband et al., 2020] with $\gamma=0.99$ and a grid size of 64x64. The figures are identical in structure to the ones in 7, 32 runs were performed for each value of $h$ in 1, 5, 10, 15, and 20. In Fig. 3, we report the number of samples to the environment at each iteration for the same runs. When compared to the figures showing the convergence, it is clearly seen that running the algorithm with higher values of lookahead results in convergence using fewer overall samples.
@@ D.4 About the Stepsizes in Simulations
In our experiments we use the exact same stepsizes described in our theory. We also observe in our simulations that the stepsize does not go to infinity but rather decreases at some point and/or stabilizes. This is a consequence of the adaptivity of the stepsize discussed above. We also note that the simulations reported in [Johnson et al., 2023] (appendix D p. 18) also validate the fact that the stepsizes do not go to infinity (see the right plots therein for the stepsizes). We also observe a similar behavior for the stepsizes in our experiments (see Figure 5).
@@ D.6 Experiments with MCTS-based Lookahead Estimation
We have performed additional experiments in an even more challenging and practical setting, on two environments from the bsuite [Osband et al., 2020] set: DeepSea and Catch. We have implemented our PMD algorithm with a Monte Carlo Tree Search (MCTS) algorithm (which we call PMD-MCTS) to compute the lookahead Q-function. We used DeepMind’s MCTS implementation (MCTX) [^3] https://github.com/google-deepmind/mctx in order to run this algorithm on any deterministic gym style environments implemented in JAX (we use gymnax [Lange, 2022] to implement these environments). Note that any tree search algorithm implementation can be used to estimate the lookahead function and plugged into our PMD method. See figures below. Note that this training procedure is more practical: we do not evaluate the value function at each state-action pair anymore at each time step, we only evaluate the value function (and so only update the policy) in states that we encounter using the current policy. This significantly relaxes the generative model assumption. Note that in this more challenging setting the best performance is not obtained for higher values of $h$: intermediate values of $h$ perform better, illustrating the tradeoff in choosing the depth $h$.
@@ D.7 Continuous Control Experiments
Finally, we performed experiments in classic continuous control environments in order to investigate the performance of $h$-PMD when applied to MDPs with large/infinite state spaces. For this continuous state space setting, we implemented a fully online variant of $h$-PMD based off of DeepMind’s MuZero [Schrittwieser et al., 2019] algorithm. Experiments show that increasing the lookahead depth can lead to better performance up to a certain point, and the tradeoff seems to be environment-specific.
Figure 8: Performance of $h$-PMD in two of OpenAI’s gym [Brockman et al., 2016] environments, CartPole-v1 (left) and Acrobot-v1 (right). In both figures the reward is plotted against the total number of environment steps evaluated, which can be understood as the number of samples used. Note that higher values of $h$ do lead to better performance for the CartPole-v1 environment, but not necessarily for the Acrobot-v1 environment.
@@ D.8 Additional experiments
We conclude with results from a larger set of experiments designed to verify the robustness of our experimental results to different environment parameters. We report the performance of the $h$-PMD algorithm in the inexact setting in two different environments, DeepSea from bsuite [Osband et al., 2020], and the classic Gridworld environment. The effect of $\gamma$ is as expected: higher values of $\gamma$ result in slower convergence, lower values of $\gamma$ yield faster convergence. The size of the environment (“chain length” for DeepSea and number of columns/rows for Gridworld) directly affects the number of iterations until convergence predictably: larger environments take more iterations to solve. Note that in all cases the effect of $h$ is as expected, i.e. higher $h$ leads to faster convergence.